A hot air balloon is rising upward with a constant speed of 20.0 m/s. When the balloon is 20.0 m above the ground, the balloonist accidentally drops a cup over the side. The balloon continue to move up at a constant speed of 20.0 m/s. Ignore the air resistance.

(1) What is the distance between the cup and balloon 1 seconds after it is dropped?

(2) How long does it take for the cup to reach the maximum height?
Answer #1

Given is:-

Initial speed of the balloon or cup is   ui = 20m / s

Height of the balloon when cup drops h = 20m

Now,

The distance covered by the cup in 1 second after it is dropped is given by second equation of motion

t2 20

by plugging all the values we get

s =(20)(1) + rac{1}{2}(-9.8)(1)^2

which gives us

s =15.1m

and the distance travelled by balloon during this 1 second is = 20m

therefore the distance between the balloon and the cup after 1 second will be

d = 20-15.1

d - 4.9m

Part-2

When the cup reach it's maximum height then it's vertical velocity will be zero therefore using first equation of motion we get

v = u + at

by plugging all the values we get

0 20 (-9.8)?

which gives us

t = 2.041

Hence,

the cup will take 2.041s to reach its maximum height after it is being dropped.